OpenStudy (anonymous):

Can this be simplified anymore? -16+4√3 over 4

OpenStudy (anonymous):

-1+\[\sqrt{3}\]

OpenStudy (precal):

\[(-16+4\sqrt{3})/4\] \[-4+\sqrt{3}\]